### v1

# Euclidean domains with no multiplicative norms

Abstract.
We construct a Euclidean domain with no multiplicative Euclidean norm to a compatibly well-ordered monoid, and hence with no multiplicative Euclidean norm to $\mathbb{R}$ (under its usual order).
A key step in the proof is showing that the UFD property is preserved when adjoining a free factorization.
Key words and phrases:
compatibly ordered monoid, Euclidean domain, Euclidean norm, unique factorization domain
2020 Mathematics Subject Classification:
Primary 13F07, Secondary 06F05, 11A05, 13A05, 13F15
1. Introduction
In this paper, we study the following well-known class of rings.
Definition 1.1.
An integral domain $R$ is called a Euclidean domain when there exists a function $\varphi\colon R-\{0\}\to\mathbb{N}$ such that for each pair $(a,b)\in R^{2}$ with $b\neq 0$ and $b\nmid a$, there exist some $q\in R$ satisfying
$\varphi(a-qb)<\varphi(b).$
Any such function $\varphi$ is called a Euclidean norm on $R$.
Readers may have seen Euclidean domains defined using a slightly different definition, perhaps with the Euclidean norm defined on 00 (just as some authors allow the zero polynomial to have a degree). As shown in [1], any such minor discrepancy causes no problems.
The historical reason for mapping to $\mathbb{N}$ is to take advantage of the fact that it possesses a well-ordering $<$, which was used to define the Euclidean norm. In the definition above, if we replace $(\mathbb{N},<)$ by an arbitrary well-ordered set $(I,<)$, we get the class of transfinite Euclidean domains, with their corresponding Euclidean norms. There are transfinite Euclidean domains that possess no $(\mathbb{N},<)$-valued Euclidean norms. In fact, for each ordinal $\alpha$, there is a transfinite Euclidean domain such that the images of its Euclidean norms must each have length larger than $\alpha$. For more information on this topic, the reader is directed to [4] and the references therein.
It has been a longstanding open problem whether or not every Euclidean domain has a multiplicative Euclidean norm, meaning that the image of $\varphi$ is in (a well-ordered subset of) a monoid $(M,\cdot,1)$ and
$\varphi(ab)=\varphi(a)\cdot\varphi(b)\ \text{ for all $a,b\in R-\{0\}$}.$
As shown recently in [4, Section 5], the answer is no if we force the monoid to be ($(\mathbb{N}-\{0\},\cdot,1)$, under its usual ordering. However, the ring constructed there does have a multiplicative Euclidean norm with values in a well-ordered submonoid of $(\mathbb{R}_{\geq 1},\cdot,1)$, under its usual ordering.
This raises the question of whether multiplicativity holds for some Euclidean norm, if we allow values in arbitrary monoids. In a technical—but somewhat trivial—sense the answer is yes, as the following proposition demonstrates.
Proposition 1.2.
Let $R$ be a Euclidean domain. There is a well-ordering $\prec$ on $R-\{0\}$ such that with respect to this ordering the identity map on $R-\{0\}$ is a (possibly transfinite) Euclidean norm for $R$.
Proof.
Let $\varphi\colon R-\{0\}\to\mathbb{N}$ be any Euclidean norm on $R$. For each $n\in\mathbb{N}$, fix a well-ordering $\prec_{n}$ on $\varphi^{-1}(n)$. Define an ordering $\prec$ on $R-\{0\}$ by saying that $x\prec y$ when either
  • (1)
    $x,y\in\varphi^{-1}(n)$, for some $n\in\mathbb{N}$, and $x\prec_{n}y$, or
  • (2)
    $x\in\varphi^{-1}(m)$ and $y\in\varphi^{-1}(n)$ for some $m,n\in\mathbb{N}$ with $m<n$.
The ordering $\prec$ is a well-ordering.
We end by showing that with respect to this ordering, the identity map on $R-\{0\}$ is a Euclidean norm. Let $a,b\in R$ with $b\neq 0$ and with $b\nmid a$. There exists some $q\in R$ with $\varphi(a-qb)<\varphi(b)$. In particular, $a-qb\prec b$. ∎
There is a significant difference between the well-ordering $\prec$ defined in the previous proposition on the multiplicative monoid $(R-\{0\},\cdot,1)$, and the well-ordering $<$ on the multiplicative monoid ($(\mathbb{N}-\{0\},\cdot,1)$. In the later case, the monoid multiplication is compatible with the ordering, in the sense that
(1.3) $a\leq b\Longrightarrow ac\leq bc\ \text{ for all $a,b,c$ in the monoid}.$
(The choice to use the nonstrict order $\leq$ when defining the compatibility condition is deliberate, as it is less stringent.) The usual ordering on $\mathbb{R}_{\geq 1}$ is also compatible with multiplication.
Do Euclidean domains always have multiplicative Euclidean norms to compatibly well-ordered monoids? We answer this important question in the negative, thus also showing that there are Euclidean domains with no $\mathbb{R}$-valued, multiplicative Euclidean norms (still assuming the values lie in a well-ordered subset, and that $\mathbb{R}$ is given its usual ordering). The remainder of the paper will be devoted to proving:
Theorem 1.4.
There exists a Euclidean domain $R$ that has no multiplicative Euclidean norm to a compatibly well-ordered monoid.
In Section 2 we state some important properties of UFDs that we will need in our construction. Readers may find Lemma 2.4 of independent interest. We end the paper by constructing an explicit ring in Section 3 that satisfies Theorem 1.4. All rings in this paper are associative, commutative, and unital. The set $\mathbb{N}$ is assumed to contain 00.
2. UFDs and splitting primes
The following well-known lemma will be used without further comment.
Lemma 2.1.
A polynomial ring in arbitrarily many variables over a UFD is a UFD.
Although adjoining polynomial variables causes no problems when working with UFDs, if we adjoin a factor of a prime to a UFD, then the new ring does not need to be a UFD. The classic example is
$\mathbb{Z}\subseteq\mathbb{Z}[\sqrt{-5}].$
Surprisingly, when adjoining new factors as freely as possible (in the universal algebra sense), there is no such problem. To motivate that result, we first discuss some notational conventions that will be used throughout this paper.
Let $R$ be a ring, let $p\in R$, and let $x$ and $y$ be two algebraically independent polynomial indeterminates. Consider the ring
$R[x,y]/(xy-p).$
This is the ring where we have freely adjoined a factorization of $p$. It is easy to see (say, by an application of the diamond lemma for ring theory [2]) that each coset in $R[x,y]/(xy-p)$ is uniquely represented by an $R$-linear combination of $1$, positive powers of $x$, and positive powers of $y$. (There are no cross terms, after repeatedly replacing any copy of $xy$ by $p$.) In particular, the natural map $R\to R[x,y]/(xy-p)$ is an embedding.
Treating $x$ as having grade $1$, treating $y$ as having grade $-1$, and treating the elements of $R$ as having grade 00, we see that $xy-p$ is homogeneous of grade 00. Thus, there is an induced $\mathbb{Z}$-grading on $R[x,y]/(xy-p)$.
Lemma 2.2.
If $R$ is an integral domain and $p\in R-\{0\}$, then $R[x,y]/(xy-p)$ is an integral domain.
Proof.
Use the standard leading terms argument, under the $\mathbb{Z}$-grading. ∎
Hereafter, we will refer to integral domains just as domains. Continue to assume that $R$ is a domain and that $p\neq 0$. Rather than handling cosets in the ring $R[x,y]/(xy-p)$, it is often more convenient to work with the isomorphic ring
$R^{\prime}:=R[s,s^{\prime}\,:\,ss^{\prime}=p].$
Throughout this paper, we will refer to $s$ and $s^{\prime}$ as conjugates. The ring $R^{\prime}$ contains $R$ as a subring. Further, $R^{\prime}$ is a domain, so it has a field of fractions. We may as well identify $s^{\prime}=p/s$. Thus, we will often write this ring as
$R^{\prime}=R[s,p/s]\subseteq R[s,s^{-1}].$
Thus, elements of $R^{\prime}$ can be written as Laurent polynomials in $s$, with coefficients from $R$. However, keep in mind that the conjugates $s$ and $s^{\prime}$ play dual roles, so we could just as easily have written
$R^{\prime}=R[s^{\prime},p/s^{\prime}]\subseteq R[s^{\prime},s^{\prime-1}].$
Since $R^{\prime}$ is a $\mathbb{Z}$-graded ring, we define the spread of a nonzero element $a\in R^{\prime}$ as the difference between the grade of the leading term of $a$ and the grade of the lowest term of $a$, writing ${\rm spread}(a)$. Thus, nonzero homogeneous elements are exactly the elements with spread 00; all other nonzero elements have positive spread. Since $R^{\prime}$ is a domain, then given $a,b\in R^{\prime}-\{0\}$ we have
(2.3) ${\rm spread}(ab)={\rm spread}(a)+{\rm spread}(b).$
The following lemma collects some additional basic facts about the ring $R^{\prime}$, under additional hypotheses.
Lemma 2.4.
Let $R$ be a UFD, and let $p$ be a prime in $R$. Put $R^{\prime}:=R[s,s^{\prime}\,:\,ss^{\prime}=p]$. The following hold:
  • (1)
    Both $s$ and $s^{\prime}$ are prime in $R^{\prime}$, and they are not associate.
  • (2)
    Any prime $q\in R$ that is not associate to $p$ remains prime $R^{\prime}$ and is not associate to either $s$ or $s^{\prime}$ in $R^{\prime}$.
  • (3)
    The units in $R^{\prime}$ are exactly the units in $R$ (i.e., $U(R^{\prime})=U(R)$).
  • (4)
    The ring $R^{\prime}$ is a UFD.
Proof.
(1) By symmetry considerations, it suffices to show that $s$ is prime in $R^{\prime}$ and not associate to $s^{\prime}$. We find that
$R^{\prime}/(s)\cong R[x,y]/(xy-p,x)\cong R[y]/(p)\cong(R/(p))[y],$
which is a domain since $p$ is prime in $R$. Further, the image of $s^{\prime}$ in this domain is nonzero (as it maps to $y$), so $s$ and $s^{\prime}$ are not associate.
(2) Let $q$ be a prime of $R$ that is not associate to $p$. We find
$R^{\prime}/(q)\cong(R/(q))[x,y]/(xy-\overline{p}).$
The element $\overline{p}\in R/(q)$ is nonzero, since $p$ is prime in $R$ and not associate to $q$. As $R/(q)$ is a domain, Lemma 2.2 shows that $(R/(q))[x,y]/(xy-\overline{p})$ is a domain. As $\overline{x}$ and $\overline{y}$ are nonzero in this ring, we see that $s$ and $s^{\prime}$ are not associate to $q$ in $R^{\prime}$.
(3) Let $a,b\in R^{\prime}$, and assume that $ab=1$. Since $R^{\prime}$ is a $\mathbb{Z}$-graded domain, we see that $a$ and $b$ must be homogeneous (of zero spread) and of opposite grades. If $a$ and $b$ both have grade 00 we are done, so it suffices to consider the case when $a$ has grade $n\in\mathbb{N}-\{0\}$, and $b$ has grade $-n$. We then can write $a=rs^{n}$ and $b=r^{\prime}s^{\prime n}$ for some nonzero elements $r,r^{\prime}\in R$. Now, $rr^{\prime}p^{n}=1$ in $R$, which is impossible since $p$ is not a unit in $R$.
(4) First, we will show that every nonzero, nonunit element $a\in R^{\prime}$ has a factorization into finitely many irreducibles. In other words, we will show that $R^{\prime}$ is an atomic domain.
In any factorization of $a$, there are at most ${\rm spread}(a)$ factors with positive spread, by (2.3\@@italiccorr). So we just need to bound the number of homogeneous, nonunit factors in any factorization. Any homogeneous factor of $a$ is also a factor of the leading term of $a$. We will show that the leading term is a finite product of primes (rather than merely irreducibles) in $R^{\prime}$, thus giving the needed bound.
By symmetry considerations, it suffices to consider the case when the leading term of $a$ is $rs^{n}$, for some $n\in\mathbb{N}$ and some $r\in R-\{0\}$. Since $R$ is a UFD, we can write
$r=uq_{1}\cdots q_{m}p^{k},$
where $u$ is a unit in $R$, the elements $q_{1},\ldots,q_{m}$ are primes in $R$ not associate to $p$ (with $m\in\mathbb{N}$), and $k\in\mathbb{N}$. Thus, up to a unit, the leading term of $a$ factors as the product $q_{1}\cdots q_{m}s^{n+k}s^{\prime k}$. These factors are all prime in $R^{\prime}$, by parts (1) and (2).
We have now shown that $R^{\prime}$ is an atomic domain. Also note that
$(s)^{-1}R^{\prime}=R[s,s^{-1}],$
which is a UFD, where $s$ is prime in $R^{\prime}$. Thus, by Nagata’s criterion (see [3, Theorem 15.41], which generalizes [5, Lemma 2]) we know that $R^{\prime}$ is a UFD. ∎
We want to iteratively use Lemma 2.4. The limit ordinal case will involve taking unions, and the following standard lemma is then useful.
Lemma 2.5.
Let $(I,<)$ be a totally ordered set, and let $(R_{i})_{i\in I}$ be a family of UFDs. Assume for all $i,j\in I$ with $i<j$ that $R_{i}\subseteq R_{j}$. Setting $R:=\bigcup_{k\in I}R_{k}$, then $U(R)=\bigcup_{k\in I}U(R_{k})$. Further, assuming that for each nonzero, nonunit element $r\in R$, there exists some index $i\in I$ (possibly depending on $r$) where
  • $r\in R_{i}$, and
  • for each index $j\geq i$, the prime factorization of $r$ in $R_{i}$ is also a prime factorization of $r$ in $R_{j}$,
then such factorizations remain prime factorizations in $R$, so $R$ is a UFD.
Proof.
Being an ordered union of domains, the set $R$ is also a domain. The claim about unit groups is also clear.
Given a nonzero, nonunit $r\in R$, fix an index $i\in I$ satisfying the two bullet points. Write the prime factorization of $r$ in $R_{i}$ as
$r=p_{1}p_{2}\cdots p_{m}.$
We will show that each $p_{i}$ is prime in $R$, thus showing that $R$ is a UFD. It suffices to show it for $p:=p_{1}$.
Given $x,y\in R$, assume that $p$ divides $xy$ in $R$. We may then write $xy=pz$ for some $z\in R$. There is then some index $j\in I$ with $j\geq i$ such that $r,x,y,z\in R_{j}$. Since $p$ is prime in $R_{j}$, then $p$ is a factor of $x$ or of $y$ in $R_{j}$. Thus, the same holds true in $R$. ∎
An iterative use of Lemma 2.4 must be done wisely, if we hope that the process will result in a UFD. For example, if we split a prime $p$ into two new primes $s_{1}$ and $s_{1}^{\prime}$, then split $s_{1}$ into two primes $s_{2}$ and $s_{2}^{\prime}$, and repeat this process recursively, then the resulting ring will not be a UFD. One must avoid infinite chains of nontrivial factors. We avoid forming such chains by never splitting a given prime more than once.
Proposition 2.6.
Let $R$ be a UFD, and let $P:=\{p_{i}\,:\,i\in I\}$ be a set of nonassociate primes in the ring $R$. Put $R^{\prime}:=R[s_{i},p_{i}/s_{i}\,:\,i\in I]$. The following hold:
  • (1)
    For each $i\in I$, both $s_{i}$ and $p_{i}/s_{i}$ are prime in $R^{\prime}$. Moreover, $s_{i}$ is not associate to $p_{i}/s_{i}$, nor is it associate to either $s_{j}$ or $p_{j}/s_{j}$, for each $j\in I$ with $j\neq i$.
  • (2)
    Any prime $q\in R$ that is not associate to a prime in $P$ remains prime in $R^{\prime}$, and it is not associate to either $s_{i}$ or $p_{i}/s_{i}$, for each $i\in I$.
  • (3)
    The units in $R^{\prime}$ are exactly the units in $R$.
  • (4)
    The ring $R^{\prime}$ is a UFD.
Proof.
We may as well replace $I$ by an ordinal $\alpha$. For each ordinal $\beta\leq\alpha$, let
$R_{\beta}:=R[s_{i},p_{i}/s_{i}\,:\,i<\beta].$
It suffices to show that each $R_{\beta}$ satisfies the four conditions above, but modified so that $R^{\prime}$ is replaced with $R_{\beta}$, the index set $I$ is replaced with $\{i\in I\,:\,i<\beta\}$, and the set $P$ is replaced with $\{p_{i}\,:\,i<\beta\}$. We will work by transfinite induction on $\beta$.
When $\beta=0$, then $R_{\beta}=R$ and all four properties hold.
Next, consider a successor ordinal $\beta+1$, and assume all four properties hold for $R_{\beta}$. Then they hold for $R_{\beta+1}$ by Lemma 2.4 (taking $R=R_{\beta}$ and $R^{\prime}=R_{\beta+1}$).
Finally, consider a limit ordinal $\beta>0$, and assume that $R_{\gamma}$ satisfies the four conditions for each $\gamma<\beta$. Note that $R_{\beta}=\bigcup_{\gamma<\beta}R_{\gamma}$. Lemma 2.5 will give us all four properties for $R_{\beta}$, once we verify that prime factorizations stabilize. Given any nonzero, nonunit element $a\in R_{\beta}$, then $a\in R_{\gamma}$ for some $\gamma<\beta$. By our inductive hypothesis, $R_{\gamma}$ is a UFD, and thus $a$ has a prime factorization in $R_{\gamma}$; write it as
$a=q_{1}q_{2}\cdots q_{m}.$
If $q_{1}$ is not associate to any $p_{i}$ with $i<\beta$, then using condition (2) and the inductive assumption, $q_{1}$ is prime in $R_{\delta}$ for each $\gamma\leq\delta<\beta$. On the other hand, if $q_{1}$ is associate to some $p_{i}$ with $i<\beta$, then $q_{1}$ factors into two primes in $R_{i}$, which remain prime in $R_{\delta}$ for each $i\leq\delta<\beta$, by the inductive assumption and condition (1). The same factorization process works for $q_{j}$, for each integer $j\in[1,m]$. Thus, the prime factorization of $a$ in $R_{\gamma}$ may factor further, but it stabilizes after a finite number of splittings. ∎
3. Main construction
We are now prepared to construct an example of a ring $R$ satisfying Theorem 1.4. Throughout this section, all notations that are introduced will continue to retain their fixed meaning when used in results.
Let $F$ be a field of characteristic zero. The assumption of characteristic zero will only be used in the proof of Proposition 3.3.
The construction consists entirely of recursively passing to a polynomial ring in multiple indeterminates, followed by passing to a subring of the field of fractions. All indeterminates are assumed to be independent of one another over $F$. The indeterminates come in three types with various subscripts. They are introduced at each successor stage $k+1$ of a recursion, with $k\in\mathbb{N}$. They are:
  • $s$-type: These are indeterminates $s_{k+1,p}$, where $p$ runs over of a certain set of primes from the $k$th stage. The $s$-indeterminates are used to split primes into products of two new prime factors.
  • $t$-type: These are indeterminates $t_{k+1}$. The $t$-indeterminates are used to give us access to indeterminates not of the other two types.
  • $u$-type: These are indeterminates $u_{k+1,a,b}$, where $a$ and $b$ run over certain elements from the $k$th stage. The $u$-indeterminates help us form certain primes that, at the very end of the construction, we localize into units. This helps guarantee that the Euclidean algorithm terminates.
To begin, let $R_{0}:=F$. View $R_{0}$ as the first step in a recursive construction of UFDs
$R_{0}\subseteq R_{1}\subseteq R_{2}\subseteq\ldots.$
We will put $R_{\infty}:=\bigcup_{k\in\mathbb{N}}R_{k}$, and the final ring $R$ that we construct will be a localization of the ring $R_{\infty}$.
Some primes in $R_{k}$ are forced to remain prime at each later stage of the recursion to prevent $R_{\infty}$ from failing to be a UFD, and they also remain prime after the final localization when pasing from $R_{\infty}$ to $R$; these are called the stable primes, the set of which is denoted $S_{k}$. Some of the primes in $R_{k}$ will split into two stable primes at the next stage; these are called the temporary primes, the set of which is denoted $T_{k}$. Finally, some primes will remain prime at each later stage of the recursion, but they will be inverted when passing from $R_{\infty}$ to $R$; these are called the unit primes, the set of which is denoted $U_{k}$.
The sets $S_{k}$, $T_{k}$, and $U_{k}$ partition the primes of $R_{k}$ (except that we allow empty parts). We will guarantee that if some prime $p\in R_{k}$ belongs to one of these three sets, then all its associates belong to the same set. At the first stage of the recursion take $S_{0}=T_{0}=U_{0}=\emptyset$ (as there are no primes in $F$).
For notational ease, once the set of temporary primes $T_{k}$ is constructed for a given integer $k\geq 0$, fix (once and for all) a subset $T_{k}^{\ast}\subseteq T_{k}$ consisting of exactly one prime from each of the association classes in $T_{k}$.
We now explain how the recursion proceeds at successor steps. Suppose that for some integer $k\geq 0$ we have been given:
  • an $F$-algebra $R_{k}$ that is a UFD, and whose units are $F^{\ast}:=F-\{0\}$, and
  • a partition of the primes of $R_{k}$ into three sets $S_{k}$, $T_{k}$, and $U_{k}$, each closed under multiplication by $F^{\ast}$.
Define $R^{\prime}_{k+1}$ to be the polynomial ring obtained by adjoining to $R_{k}$ the following three types of indeterminates:
  • (1)
    $s_{k+1,p}$, for each $p\in T_{k}^{\ast}$,
  • (2)
    $t_{k+1}$, and
  • (3)
    $u_{k+1,a,b}$, for each pair of elements $a,b\in R_{k}-\{0\}$, where $\gcd_{R_{k}}(a,b)=1$, and $b$ is not divisible (in $R_{k}$) by any element of $T_{k}$, nor by the square of any element in $S_{k}$.
Hereafter, subscripts on these variables will automatically be assumed to belong to the appropriate sets, subject to the restrictions above.
Finally, fix
$R_{k+1}:=R^{\prime}_{k+1}[p/s_{k+1,p}\,:\,p\in T_{k}^{\ast}],$
which is a subring of the fraction field of $R^{\prime}_{k+1}$. In fact, by Proposition 2.6, $R_{k+1}$ is a UFD containing $R_{k}$, whose units are still exactly $F^{\ast}$. Moreover, the stable primes in $S_{k}$ and the unit primes in $U_{k}$ all stay prime in $R_{k+1}$, and each of the temporary primes in $T_{k}$ has split into two new primes.
For notational convenience, we take $s^{\prime}_{k+1,p}:=p/s_{k+1,p}$, which is the conjugate of $s_{k+1,p}$. Now, fix the new set of stable primes to be
$S_{k+1}:=S_{k}\cup(F^{\ast}\cdot\{s_{k+1,p},s^{\prime}_{k+1,p}\,:\,p\in T_{k}^{\ast}\}).$
Also, fix the new set of unit primes to be
$U_{k+1}:=U_{k}\cup(F^{\ast}\cdot\{a-u_{k+1,a,b}b\,:\,u_{k+1,a,b}\in R_{k+1}^{\prime}\}),$
Note that $a-u_{k+1,a,b}b$ really is prime in $R_{k+1}$, since this is a linear polynomial in the indeterminate $u_{k+1,a,b}$, and the two coefficients are nonzero and relatively prime (in $R_{k}$, and hence also in $R_{k+1}$). Clearly, the unit primes are not $F^{\ast}$ multiples of the stable primes. (The unit primes will become units at the very end of the construction.) Notice that with these definitions, we have two nondecreasing chains
$S_{0}\subseteq S_{1}\subseteq S_{2}\subseteq\ldots$
and
$U_{0}\subseteq U_{1}\subseteq U_{2}\subseteq\ldots$
of sets of primes. Finally, fix $T_{k+1}$ to be the set of all remaining primes in $R_{k+1}$ (and fix a transversal $T_{k+1}^{\ast}$ of the association classes in $T_{k+1}$).
This finishes the recursion. We now take
$R_{\infty}:=\bigcup_{k\in\mathbb{N}}R_{k},\qquad S_{\infty}:=\bigcup_{k\in\mathbb{N}}S_{k},\qquad U_{\infty}:=\bigcup_{k\in\mathbb{N}}U_{k}.$
By Lemma 2.5, the ring $R_{\infty}$ is a UFD, the set of whose primes is $S_{\infty}\cup U_{\infty}$, which we continue to call the stable primes and unit primes, respectively.
Note that $R_{\infty}$ is a subring of the Laurent polynomial ring in the $s$-type indeterminates, with coefficients that are polynomials over $F$ in the other indeterminates. Viewed this way, any element $r\in R_{\infty}$ can be written in the unique form $c/d$, where $d$ is a (finite, possibly empty) product of $s$-type indeterminates, $c$ is an $F$-linear combination of distinct monomials in the indeterminates (with nonzero coefficients), and no $s$-type variable that occurs in $d$ also occurs in all the monomials in the support of $c$.
Definition 3.1.
Given any $r\in R_{\infty}$, the rank of $r$ is the smallest integer $k\geq 0$ such that $r\in R_{k}$, denoted by $\operatorname{rank}(r)$.
The rank of any indeterminate is its first subscript. Given any $r\in R_{\infty}-F$, then its rank is easy to determine; it is the largest of all the ranks of the indeterminates that appear in its reduced form. Given a unit prime $a-u_{k+1,a,b}b$, note that $u_{k+1,a,b}$ is the unique indeterminate of maximal rank that appears in (at least one monomial in the support of) the prime; we will call $u_{k+1,a,b}$ the ranking indeterminate of the unit prime. The following fact about products of unit primes is easy, but we include the proof for completeness.
Lemma 3.2.
The rank of a nonempty product of unit primes is the maximum of the ranks of those primes.
Proof.
Consider any nonempty product $\prod_{i=1}^{n}a_{i}$ of nonzero elements from $R_{\infty}$, for some integer $n\geq 1$. Write the reduced form for $a_{i}$ as $c_{i}/d_{i}$. Once $c:=\prod_{i=1}^{n}c_{i}$ is expanded, it is an $F$-linear combination of distinct monomials (with nonzero coefficients), where the indeterminates in those monomials also appeared in $c_{1},c_{2},\ldots,c_{n}$. Moreover, any indeterminate appearing in some $c_{i}$ must still appear in at least one monomial in the support of $c$. Note that $d=\prod_{i=1}^{\infty}d_{i}$ is a monomial in $s$-type variables.
The only thing that would prevent $c/d$ from being the reduced form for the product is that some $s$-type variable in the denominator could appear in each of the monomials in the support of the numerator, and cancel off. The other two types of indeterminates do not cancel off, and in particular if each $a_{i}$ is a unit prime, the ranking indeterminates all appear in the reduced form. The largest rank of those ranking indeterminates clearly bounds the ranks of all other indeterminates in the expression $c/d$. ∎
The previous lemma applies to more general situations. However, note that it is false when applied to stable primes instead of unit prime, for if $s_{k+1,p}$ is any $s$-type variable, then $p$ factors as $s_{k+1,p}\cdot(p/s_{k+1,p})$, but $p$ has smaller rank than either of the factors. Of course, the rank of a product can never be bigger than the maximum of the ranks of the factors.
Another example of using rank considerations is as follows. Let $r\in R_{\infty}-F$ have rank $k\geq 1$. Let $x$ be a $t$-type or $u$-type indeterminate of rank $k$. Write $r$ in its unique form $c/d$. Treating $c$ as a polynomial in the indeterminate $x$, we can then write $c=c_{0}+c_{1}x+\cdots+c_{n}x^{n}$ for some integer $n\geq 0$, where each $c_{i}$ is a polynomial in the other indeterminates that appear in $c$. We claim that $c_{i}/d\in R_{\infty}$ for each integer $i\in[0,n]$. To see this, note that in the recursive definition of $R_{k}$ we could just as easily have adjoined the single indeterminate $x$ after all the other adjunctions, and so indeed any element of $R_{k}$ is a polynomial in the variable $x$. When $x$ is an $s$-type indeterminate, we can treat $r$ as a Laurent polynomial in $x$ and again conclude that each coefficient belongs to $R_{\infty}$ (by essentially the same argument).
Note that the result of the previous paragraph may fail when $x$ has rank smaller than $r$. For example, $1-t_{1}\in R_{1}$ is a temporary prime in $R_{1}$, and so $(1-t_{1})/s_{2,1-t_{1}}$ is an element of $R_{2}$. Writing this element as a polynomial in the variable $x:=t_{1}$, we have
$\frac{1}{s_{2,1-t_{1}}}+\frac{-1}{s_{2,1-t_{1}}}t_{1},$
but the coefficient $1/s_{2,1-t_{1}}$ does not belong to $R_{2}$ (or even $R_{\infty}$).
We are now ready to prove the following key property of the ring $R_{\infty}$.
Proposition 3.3.
Given any $p,q,v\in R_{\infty}$ with $p\in S_{\infty}$ and $v\in F^{\ast}\cdot\langle U_{\infty}\rangle$, then
$vt_{\operatorname{rank}(p)+1}-qp^{2}$
has at least two prime factors from $S_{\infty}$, counting multiplicity.
Proof.
Fix $k:=\operatorname{rank}(p)$, and fix $r:=vt_{k+1}-qp^{2}$. After removing any common (unit prime) factors of $v$ and $q$, we may as well assume $\gcd_{R_{\infty}}(q,v)=1$. There are three main cases.
Case 1: $k+1\geq\max(\operatorname{rank}(v),\operatorname{rank}(q))$. This means that $t_{k+1}$ cannot appear in any prime factor of $v$ (in $R_{\infty}$), because the corresponding ranking indeterminate would then have rank greater than $k+1$, making $\operatorname{rank}(v)>k+1$ by Lemma 3.2.
Next, assume by way of contradiction that $t_{k+1}$ does not appear in the reduced form for $r$. As the reduced form for $vt_{k+1}$ is linear in the indeterminate $t_{k+1}$, the same must be true for $q$. As $q\in R_{k+1}$, we can then write $q=q_{0}+q_{1}t_{k+1}$, where $q_{0},q_{1}\in R_{k+1}$ do not have $t_{k+1}$ in their reduced forms. Now, since $t_{k+1}$ does not appear in $r$, we must have $v=q_{1}p^{2}$. Then $p\mid v$, which is impossible (as $v$ has only unit prime factors).
So, $t_{k+1}$ does appear in the reduced form for $r$. Therefore, it must appear in one of the prime factors of $r$ in $R_{k+1}$. It cannot appear in a stable prime (since that would involve an $s$-type variable of larger rank, hence the factor wouldn’t belong to $R_{k+1}$). It cannot appear in a unit prime (since that would involve a $u$-type variable of larger rank). Thus, it belongs to a temporary prime, which factors into two stable primes at the next stage, and so $r$ has at least two stable prime factors in $R_{\infty}$.
Case 2: $k+1<\operatorname{rank}(v)$ and $\operatorname{rank}(q)\leq\operatorname{rank}(v)$. Fix $\ell:=\operatorname{rank}(v)$. By Lemma 3.2, there is a unit prime factor $a-u_{\ell,a,b}b$ of $v$. Fix $u:=u_{\ell,a,b}$, and write $v=(a-ub)^{\alpha}v^{\prime}$, for some integer $\alpha\geq 1$ and some $v^{\prime}\in R_{\infty}$ with $(a-ub)\nmid v^{\prime}$. (In other words, $\alpha$ is the $(a-ub)$-adic valuation of $v$.) Notice that $u$ does not appear in $v^{\prime}$, by rank considerations.
Next, assume by way of contradiction that $u$ does not appear in the reduced form for $r$. There are two subcases to consider. First, we may have $p\mid b$. Notice that from the definition of the $u$-type variables, this means that $p\nmid a$. Now, treating $v$ as a polynomial in the variable $u$, the coefficient of the linear term is $-\alpha a^{\alpha-1}bv^{\prime}$. On the other hand, as $q\in R_{\ell}$, we can write
$q=q_{0}+q_{1}u+\text{ higher order terms}$
where $q_{0},q_{1},\ldots\in R_{\ell}$, and $u$ does not appear in the reduced form of any $q_{i}$. Thus, since $u$ does not appear in the reduced form for $r$, we must have $p^{2}q_{1}=-\alpha a^{\alpha-1}bv^{\prime}t_{k+1}$. As we are working over a characteristic zero field, $\alpha\neq 0$. Hence, $p^{2}$ divides $b$ (since $p\nmid av^{\prime}t_{k+1}$). This contradicts the fact that the third subscript on a $u$-type variable cannot be divisible by the square of a stable prime.
The second subcase is when $p\nmid b$. Now, looking at the coefficient of $u^{\alpha}$, rather than the linear coefficient, we obtain a similar contradiction.
Thus, we now know that $u$ does appear in the reduced form for $r$, hence in one of its prime factors in $R_{\ell}$. It cannot appear in a stable prime, or in a unit prime where $u$ is not the ranking indeterminate (by considering ranks, as in the last paragraph of Case 1). Also, $a-ub$ cannot divide $r$, else it divides $q$, contradicting the fact that $\gcd_{R_{\infty}}(q,v)=1$. Thus, $u$ appears in a temporary prime, and we are done as in Case 1.
Case 3: $k+1<\operatorname{rank}(q)$ and $\operatorname{rank}(v)<\operatorname{rank}(q)$. Fix $\ell:=\operatorname{rank}(q)$. If there is a $t$-type variable of this rank in $q$, then we finish as in the last paragraph of Case 1. If there is a variable $u:=u_{\ell,a,b}$ of this rank in $q$, then we finish as in the last paragraph of Case 2, except when $r=(a-ub)^{\alpha}r^{\prime}$, where $r^{\prime}\in R_{\ell}$ does not have $u$ in its reduced form. (We must consider this possibility, since we do not have $(a-ub)\mid v$.) If $p\nmid b$, then considering the coefficient of $u^{\alpha}$ we must have $p\mid r^{\prime}$. But then $p\mid r$, and so $p\mid vt_{k+1}$, which is impossible. If $p\mid b$, then considering the coefficient of $u$, we must have $p\mid r^{\prime}$, leading to the same contradiction.
Thus, we may reduce to the case that there are no $t$-type nor $u$-type variables of rank $\ell$ in $q$. Therefore, there is a variable $s:=s_{\ell,p^{\prime}}$ that must appear in the reduced form for $q$, and hence for $r$. It must then also appear in at least one of the prime factors of $r$ in $R_{\ell}$. If such a factor is a temporary prime, we are done as before. If such a factor is a unit prime, this contradicts the fact that we are working in $R_{\ell}$ since the ranking variable would have rank bigger than $\ell$. So, we reduce to the case that any such prime factor must be $s$ or its conjugate. If there are two or more such factors, we are done. So we may assume (by way of contradiction) that there is just one such prime factor. By symmetry considerations, it suffices to consider the case when that single prime factor is $s$.
Writing $q$ as a Laurent polynomial in the indeterminate $s$, then it is of the form
$q=s^{-m}(q_{0}+q_{1}s+\cdots+q_{n}s^{n}),$
for some integers $m,n\geq 0$, where each $q_{i}\in R_{\ell}$ does not have $s$ appear in its reduced form (for each integer $i\in[0,n]$), and where $q_{n}$ is nonzero and $n$ is minimal (so $m$ is also minimal). If $m>0$, then the reduced form for $r=vt_{\operatorname{rank}(p)}-qp^{2}$ cannot be linear in $s$, as described in the previous paragraph. Thus $m=0$, and a similar argument shows that $n=1$. This forces $vt_{\operatorname{rank}(p)+1}=q_{0}p^{2}$, which is impossible since $p$ does not divide the left side. ∎
As already mentioned, $R_{\infty}$ is a UFD. We finally invert the unit primes and consider the new ring $R:=U_{\infty}^{-1}R_{\infty}$. This is a UFD where the primes of $R$ are exactly the associates (in $R$) of the elements in $S_{\infty}$ (i.e., the primes in $R_{\infty}$ that were not inverted).
Corollary 3.4.
The ring $R$ does not possess a multiplicative Euclidean norm to a compatibly well-ordered monoid.
Proof.
Assume, by way of contradiction, that there is a compatibly well-ordered monoid $(M,\cdot,1_{M},\leq)$ and a multiplicative Euclidean norm $\psi\colon R-\{0\}\to M$ (with respect to the strict version of this ordering).
Note that $1_{M}\in M$ is the smallest element of $M$, for if $1_{M}>m$, then multiplying by $m$ we get $m>m^{2}$, and repeatedly multiplying by $m$ produces the infinite descending chain
$1_{M}>m>m^{2}>\cdots,$
contradicting the fact that $\leq$ is a well-ordering. In particular, $\psi(r)\geq 1_{M}$ for each $r\in R-\{0\}$.
Fix a prime $p\in R$ such that $\psi(p)$ is as small as possible. Now, since $\psi$ is a Euclidean norm, there exists some $q^{\prime}\in R$ with
$\psi(t_{\operatorname{rank}(p)+1}-q^{\prime}p^{2})<\psi(p^{2}).$
We can write $q^{\prime}=v^{-1}q$ for some $q\in R_{\infty}$ and some $v\in F^{\ast}\cdot\langle U_{\infty}\rangle$. Since $\psi$ is invariant on unit multiples, we have
$\psi(vt_{\operatorname{rank}(p)+1}-qp^{2})=\psi(t_{\operatorname{rank}(p)+1}-q^{\prime}p^{2}).$
By Proposition 3.3, $(vt_{\operatorname{rank}(p)+1}-qp^{2})=p_{1}p_{2}r$ for some stable primes $p_{1},p_{2}\in S_{\infty}$ and some element $r\in R-\{0\}$. Using the multiplicative nature of $\psi$, and the minimality property of $\psi(p)$, we find the contradiction
$\psi(p^{2})=\psi(p)\psi(p)\leq\psi(p_{1})\psi(p_{2})1_{M}\leq\psi(p_{1})\psi(p_{2})\psi(r)=\psi(vt_{\operatorname{rank}(p)+1}-qp^{2})<\psi(p^{2}).$
(Note that we implicitly used the compatibility condition (1.3\@@italiccorr), in order to obtain both of the nonstrict inequalities.) ∎
We finish proving Theorem 1.4 by showing:
Theorem 3.5.
The ring $R$ is a Euclidean domain.
Proof.
Define a function $\varphi\colon R-\{0\}\to\mathbb{N}$ by the rule
$u\prod_{i=1}^{k}p_{i}^{\alpha_{i}}\mapsto\sum_{i=1}^{k}\alpha_{i}^{2},$
where the $p_{i}$ are nonassociate primes in $R$, the $\alpha_{i}$ are nonnegative integers, and where $u$ is a unit. Note that $\varphi$ is invariant when multiplying its argument by a unit, but its value increases after multiplying by a (nonzero) nonunit.
We will show that $\varphi$ is a Euclidean norm on $R$. To that end, let $a,b\in R$ with $b\neq 0$. Write $a=ga^{\prime}$ and $b=gb^{\prime}$, with $a^{\prime},b^{\prime}\in R_{\infty}$ and $g\in R$, and with $\gcd_{R_{\infty}}(a^{\prime},b^{\prime})=1$. (In other words, take $g:=\gcd_{R}(a,b)$, but since GCDs are only defined up to unit multiples, we also guarantee that the cofactors belong to $R_{\infty}$.) We may then fix an integer $k\geq 0$ such that $a^{\prime},b^{\prime}\in R_{k}$. Note that we can also guarantee that $\gcd_{R_{k}}(a^{\prime},b^{\prime})=1$, after absorbing any common unit primes into $g$. Further, after increasing $k$ by $1$ if necessary, we may assume that all the prime factors of $a^{\prime}$ and $b^{\prime}$ from $R$ already live in $R_{k}$. There are three cases to consider.
Case 1: $b^{\prime}\mid a^{\prime}$ in $R$. In this case $b\mid a$, and we are done. For all remaining cases, we will assume $b^{\prime}\nmid a^{\prime}$. In particular, $b^{\prime}$ is not a unit in $R$ and $a^{\prime}\neq 0$.
Case 2: $b^{\prime}$ has no repeated prime factor in $R$. Taking $q:=u_{k+1,a^{\prime},b^{\prime}}$, then $a^{\prime}-qb^{\prime}$ is a unit of $R$. Thus, since $b^{\prime}$ is not a unit in $R$,
$\varphi(a-qb)=\varphi((a^{\prime}-qb^{\prime})g)=\varphi(g)<\varphi(gb^{\prime})=\varphi(b),$
as desired.
Case 3: $b^{\prime}$ has a repeated prime factor in $R$ (so $\varphi(b^{\prime})\geq 4>2$). Taking
$\ell:=\max(\operatorname{rank}(a^{\prime}),\operatorname{rank}(b^{\prime}),\operatorname{rank}(g))$
and putting $q:=t_{\ell+1}$, then $a^{\prime}-qb^{\prime}$ is a temporary prime in $R_{\ell+1}$, and hence it factors as a product of two new, conjugate stable primes in $R_{\ell+2}$, which are thus coprime to $g$ by rank considerations. Therefore,
$\varphi(a-qb)=\varphi((a^{\prime}-qb^{\prime})g)=2+\varphi(g)<\varphi(b^{\prime})+\varphi(g)\leq\varphi(b^{\prime}g)=\varphi(b),$
where the last (nonstrict) inequality comes from the fact that $\alpha^{2}+\beta^{2}\leq(\alpha+\beta)^{2}$, for all integers $\alpha,\beta\geq 0$. ∎
Interestingly, this proof also shows that there is a Euclidean algorithm for $R$ that terminates with a zero remainder after at most three steps. Also note that $R$ has the same cardinality as $F$, which may be any infinite cardinality since $F$ has characteristic zero.
4. Acknowledgements
We thank Jesse Elliott for pointing out an interesting reference, Luc Guyot for comments that improved the proof of Lemma 2.4, and Kyle Pratt for many comments that improved the quality of this paper. This work was partially supported by a grant from the Simons Foundation (#963435 to Pace P. Nielsen).